# A Framework for Double-Blind Federated Adaptation of Foundation Models

### v1

@@ B.3 Approximation Results
Softmax.
Given a vector ${\bm{z}}\in\mathbb{R}^{n}$ with components $z_{i}$, the softmax function is defined as:
@@
for some $\xi$ between 00 and $x$, indicating the error in approximating $e^{x}$ with $P_{d}(x)$.
Softmax approximation results. Table 5 presents softmax approximations while keeping all other non-linear components fixed. Using Eq. 2, we vary the polynomial degree $d$ for the approximation during a single epoch of knowledge distillation (Section A, Stage I) and compare the results to the true softmax. The table reports key performance metrics, including attention loss $\mathcal{L}_{a}$ (Eq. 12), representation loss $\mathcal{L}_{h}$ (Eq. 13), and accuracy. Additionally, it provides efficiency metrics such as the GPU memory usage and latency for processing a batch of samples. Balancing the trade-offs between time, memory requirements, and performance drop, we chose $d=6$ for all experiments.
Degree $d$ Memory (GB) Latency (ms) Attention loss $(\mathcal{L}_{a})$ Representation loss $(\mathcal{L}_{h})$ One-epoch accuracy

### v2

@@ D.2 Main Results
D.3 Distillation Results
Following the findings discussed in Section C, we present performance plots for two key processes: (i) fine-tuning the teacher model $\mathcal{T}$ on a public auxiliary dataset $\mathcal{D}_{aux}$, and (ii) distilling the student model $\mathcal{S}$ on the same dataset $\mathcal{D}_{aux}$ using the trained teacher model $\mathcal{T}$. Figure 14 illustrates the performance metrics - balanced accuracy and loss - of the fine-tuning process on the Tiny-ImageNet dataset. After completing this training, we proceed with distillation for the approximation-enabled student model. Figure 14 shows the performance plot for the distillation process on the Tiny-ImageNet dataset. In the loss behavior plot (on the right), a dashed line marks the start of the second stage. During the first stage, as described in Eq. 21, transformer-layer distillation is performed (showing a higher value due to the high-dimensional attention and hidden representation matrices). Then, the second-stage distillation follows, which is a prediction-layer distillation [19]. The learning rate scheduler is employed at epochs $15$ and $25$ with a decay factor of $0.1$. Figures 14 and 14 present analogous performance plots for the Fed-ISIC2019 dataset, using center=0 as the public auxiliary dataset $\mathcal{D}_{aux}$. The learning rate scheduler is employed at epochs $20$ and $40$ with a decay factor of $0.1$.
D.4 Approximation Results
Softmax. Given a vector ${\bm{z}}\in\mathbb{R}^{n}$ with components $z_{i}$, the softmax function is defined as:
@@ D.4 Approximation Results
for some $\xi$ between 0 and $x$, indicating the error in approximating $e^{x}$ with $P_{d}(x)$.
Degree $d$ Memory (GB) Latency (ms) Attention loss $(\mathcal{L}_{a})$ Representation loss $(\mathcal{L}_{h})$ One-epoch accuracy $(\%)$
@@
Softmax approximation results. Table 5 presents softmax approximations while keeping all other non-linear components fixed. Using Eq. 3, we vary the polynomial degree $d$ for the approximation during a single epoch of knowledge distillation (Section C, Stage I) and compare the results to the true softmax. The table reports key performance metrics, including attention loss $\mathcal{L}_{a}$ (Eq. 17), representation loss $\mathcal{L}_{h}$ (Eq. 18), and accuracy. Additionally, it provides efficiency metrics such as GPU memory usage and latency for processing a batch of samples. Balancing the trade-offs between time, memory requirements, and performance drop, we chose $d=6$ for all experiments.
@@
Time taken to encrypt one sample $1062$ ms
Time taken to decrypt one sample $168.7$ ms
Time taken to encrypt a batch of 8 samples $4.65$ s
Time taken to decrypt a batch of 8 samples $231.59$ ms
Ciphertext size of one sample $17.33$ MB
Plaintext size of one sample $6.21$ MB
Encrypted inference time $136$ s
Table 7: Computational and memory overhead for encrypted inference and encryption using FHE with the Tile Tensors framework. Results are presented per transformer block, evaluated at a 128-bit security level, including approximations for Softmax, inverse, and GELU operations.
D.5 Homomorphic Encryption Results
For the implementation of fully homomorphic encryption (FHE), we utilize the Tile Tensors framework [2, 8] on an NVIDIA A100-PCIE-40GB machine with 250 GB of available RAM. Table 7 summarizes the computational and memory requirements for our experiments, all evaluated at the standard 128-bit security level. The results presented in the table correspond to the encryption of one transformer block, accounting for approximations of operations such as Softmax, inverse, and GELU. Specifically, we report both the encrypted inference time and ciphertext memory size. The encrypted inference is executed on a powerful server, and the required time determines its computational complexity and latency, and the ciphertext size directly determines the communication overhead within the federated setting.
The encryption of a single data sample with a dimension $(577,768)$ takes approximately $1062$ ms, while decrypting the same sample takes significantly less time, around $168.7$ ms, highlighting the asymmetry in encryption and decryption costs. Batch processing improves efficiency, as encrypting $8$ samples takes $4.65$ seconds, whereas decryption for the same sized batch completes in $231.59$ ms. The ciphertext size of a single encrypted sample expands to $17.33$ MB, larger than its plaintext counterpart ($6.21$ MB), indicating the additional communication overhead introduced by encryption. Moreover, encrypted inference for one transformer block requires approximately $136$ seconds on the specified hardware, highlighting the computational overhead of inference under encryption. Nonetheless, this inference time could be substantially reduced with optimized cryptographic primitives, more powerful server infrastructure, and GPU acceleration [27, 25, 60, 44].
Computational Overhead. As with any cryptographic approach for privacy-preserving ML, BlindFed introduces non-negligible computational costs. However, it is designed to minimize the burden on thin clients by offloading most computations to a powerful server (Table 7). On the client side, only encryption/decryption of intermediate representations and a plaintext parallel adapter update are performed. On commodity CPUs, encrypting (decrypting) a single sample takes $\mathbf{1.06\,s}$ ($\mathbf{0.17\,s}$), and encrypting (decrypting) $8$ samples in batch takes $\mathbf{4.7\,s}$ ($\mathbf{0.23\,s}$), requiring less than $\mathbf{1\,GB}$ of RAM. The server handles the expensive encrypted inference for each ViT block under CKKS, taking roughly $\mathbf{136\,s}$ per sample (amortized) and requiring more than $\mathbf{22\,GB}$ of memory. However, as mentioned above, these operations are embarrassingly parallel across clients, and GPU-accelerated CKKS implementations can further reduce the server-side inference time by over $10\times$.